OpenStudy (anonymous):

simplify 4 radical 2 + radical 32 - radical 16

OpenStudy (anonymous):

\[4\sqrt{2}+8\sqrt{2}-4\sqrt{2}\]

OpenStudy (anonymous):

Now you basically just add and subtract

OpenStudy (anonymous):

JL look at radical 32 gain, you've done a mistake there

OpenStudy (anonymous):

Are you sure? I'm pretty sure that's right

OpenStudy (anonymous):

\[\sqrt{32} = \sqrt{16*2} = 4\sqrt{2}\]

OpenStudy (anonymous):

also radical 16 you've done the same mistake, just re check

OpenStudy (anonymous):

Oops...guess I wasn't paying attention so its \[4\sqrt{2}+4\sqrt{2}+4\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but what do you do with the like terms?